Circular No. 8170

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 O1 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to have cometary appearance on CCD images taken by P. Kušnirak (Ondřejov; 0.65-m f/3.6 reflector; well-condensed condensation and a faint 20′′ tail toward the southeast) and by P. Birtwhistle (Great Shefford, U.K.; nuclear condensation of diameter ∼ 6′′ with a faint, short, broad tail ∼ 15′′ long in p.a. 139^{o}; mag 17.3–18.2).

2003 UT  α_{2000}  δ_{2000}  Mag.
July 19.26109  19^{h}34^{m}24.^{s}63  +38^{o}36′12′′.4  18.4

Additional astrometry is given on MPECs 2003-O15 and 2003-O40; the following parabolic orbital elements and an ephemeris also appear on MPEC 2003-O40.

T = 2004 Mar. 10.367 TT  ω = 81.^{o}104
Ω = 347.607 }2000.0
q = 6.85733 AU  i = 117.967

SUPERNOVAE 2003gn, 2003gp, AND 2003gq A. V. Filippenko, R. J. Foley, and L. Desroches, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 320–1000 nm) obtained on July 27 UT with the Shane 3-m telescope at Lick Observatory reveals that SN 2003gn (IAUC 8168) is of type Ia, close to maximum brightness. SN 2003gq (IAUC 8168) is also of type Ia, but its spectrum resembles that of the peculiar SN 1991T (e.g., Filippenko et al. 1992, Ap.J. 384, L15) near or slightly before maximum brightness: the Si II and Ca II absorption lines are weak, while Fe III lines are prominent. SN 2003gp (IAUC 8168) is of type II, not long after the explosion; the continuum is quite blue, and low-contrast Hα having a P-Cyg profile is visible.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003gn, obtained by P. Berlind on July 29.37 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 10328 km/s for the host galaxy, the supernova expansion velocity is 12800 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 1 ± 2 days before maximum light.

2003 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green